OpenStudy (anonymous):

Solve the quadratic equation by completing the square. x^2-12x+34=0 First, choose the appropriate form and fill in the blanks with the correct numbers. Then, solve the equation. Round your answer to the nearest hundredth. If there is more than one solution, separate them with commas.

OpenStudy (freckles):

\[x^2-12x=-34 \\ x^2-12x+?=-34+? \] recall \[x^2+kx+(\frac{k}{2})^2=(x+\frac{k}{2})^2\]

OpenStudy (freckles):

so what do we need to add to both sides in order to right the left hand side as something square (hint: first step notice the coefficient of x)